Tohban Report 2013-04-03

From RHESSI Wiki

Jump to: navigation, search

Tohban Reports
Start Date: 27 Mar 2013
End Date: 04 Apr 2013
Tohban: Mitsuo Oka
Tohban email:
Next Tohban: Juan Carlos Martinez Oliveros
List all reports


Solar Activity

Solar activity remained low with only 2 C-class flares. These C-class flares occurred on the last day of the week (Apr-03) and solar activity may become moderate over the next few days. On the other hand, STEREO B shows no active region which will rotate onto the disk over the next few days.

How many GOES flares occurred?

 Flares above B, C, M, X class were     11     0     0     0

And how many of these are listed in the RHESSI flare list?

 Flares above B, C, M, X class were      7     0     0     0

And how many had EXCELLENT coverage?

 Flares above B, C, M, X class were      3     0     0     0

There were RHESSI flares/GOES flares 124 / 11 over the time range 26-Mar-13 02-Apr-13

Memory Management

The SSR emptied after most pass set.

Spacecraft Status

A spin-up from 14.6 to 15.0 RPM was performed on 2013-086 between 18:23 and 19:30 UT.

Data Gaps

There is one small (300 sec) data gap in VC3 data. GAP START TIME GAP END TIME GAP (SEC) 2013-04-02T23:55:00.000 -- 2013-04-03T00:00:00.000 300.00000

Detector issues

Detector 5 HV was lowered by 10 counts, bringing the voltage to 3800 V. The detector 5 monitor rate remains elevated (as high as that of detector 2).

Other notes

The problem of quicklook generation processing has been fixed.

Decimation Normal/vigorous
HLAT Decimation Rear decimation weight 6, no front decimation
Night time data (fronts) +/- 4 minutes
Night time data (rears) Taking data at nights
Require extra passes? No
Requirement for moving pointer? No
Attenuator operation Normal
Detector problems? Detector 5 arc + extended noise period.
Personal tools